Determine the second derivative of g(x)=sin(2x³-9x)


1
Expert's answer
2022-05-18T09:32:06-0400
"g'(x)=(\sin(2x^3-9x))'=(6x^2-9)\cos (2x^3-9x)"

"g''(x)=((6x^2-9)\cos (2x^3-9x))'"

"=12x\cos (2x^3-9x)-(6x^2-9)^2\sin (2x^3-9x)"